# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v1

@@ 3.3 Latent Variable Framework: Energy Function and Inference
    Initialize $\bm{z}^{(0)}$ (e.g., a zero vector or from the Encoder).
@@
    For $k=0$ to $K-1$:
    • Compute the (sub)gradient of the energy $E$ with respect to $\bm{z}$ at $\bm{z}^{(k)}$, holding $\bm{h}$ and $\theta_{\text{Dec}}$ constant [14]:
      $\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})=\frac{\partial C}{\partial\text{Dec}}\cdot\frac{\partial\text{Dec}(\bm{z}^{(k)},\bm{h})}{\partial\bm{z}}+\lambda\,\text{sign}(\bm{z}^{(k)})$
    • Update $\bm{z}$ using a step size $\alpha_{\text{inf}}$:
      $\bm{z}^{(k+1)}\leftarrow\bm{z}^{(k)}-\alpha_{\text{inf}}\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})$
@@
    The result $\bm{z}^{(K)}$ serves as the inferred optimal latent variable for this sample: $\bm{z}^{*}\approx\bm{z}^{(K)}$.
This inference process finds a sparse latent code $\bm{z}^{*}$ under the current model parameters that aims to best explain the observed future $Y$ given the history $\bm{h}$.
3.4 Amortized Inference via Encoder (Enc)
The $K$-step inference process, while crucial for finding an optimal latent variable $\bm{z}^{*}$, introduces computational overhead during training and especially at test time if performed for every sample. To address this, the Encoder module $\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$ provides amortized inference by learning to directly map the observed future $Y$ and historical context $\bm{h}$ to an estimate $\hat{\bm{z}}$ of $\bm{z}^{*}$ [33].
$\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
This learned mapping enables faster inference once the model is trained. During training, the explicit $K$-step inference process is used to generate the target $\bm{z}^{*}$ for supervising the Encoder’s learning. At test time, for forecasting, the Encoder cannot use the future $Y$; instead, an alternative encoder $\text{Enc}_{\text{test}}(X;\theta_{\text{EncTest}})$ might be trained to predict $\bm{z}^{*}$ from $X$ only, or the $K$-step inference is run starting from $\bm{z}^{(0)}=0$. This paper uses the latter approach for generating $\bm{z}^{*}$ at test time.
3.5 Training Objective and Parameter Updates
The training process aims to minimize errors in both prediction (using the inferred latent variable $\bm{z}^{*}$) and in the amortized inference of $\bm{z}^{*}$ by the Encoder (if an Encoder for $\bm{z}^{*}$ from $X$ and $Y$ is used for training efficiency). The parameters $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{Enc}}$ are updated based on two primary loss components, calculated after obtaining $\bm{h}=\text{Pred}(X)$, inferring $\bm{z}^{*}$ via $K$-step optimization, and, if applicable, computing $\hat{\bm{z}}=\text{Enc}(Y,\bm{h})$:
  1. 1.
    Prediction Loss ($L_{\text{Pred}}^{*}$): Measures the error of the Decoder when using the inferred optimal latent variable $\bm{z}^{*}$.
    $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$
  2. 2.
    Inference Loss ($L_{\text{Inf}}$) (Optional, for training the Encoder Enc): Measures the discrepancy between the optimal latent variable $\bm{z}^{*}$ (from K-step inference) and the training-time Encoder’s prediction $\hat{\bm{z}}$.
    $L_{\text{Inf}}=D(\bm{z}^{*},\hat{\bm{z}})=\left\|\bm{z}^{*}-\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})\right\|_{2}^{2}$
The gradients for updating the network parameters are derived from these losses:
  • $\theta_{\text{Dec}}$ is updated based on $\nabla_{\theta_{\text{Dec}}}L_{\text{Pred}}^{*}$.
  • $\theta_{\text{Enc}}$ (if used) is updated based on $\nabla_{\theta_{\text{Enc}}}L_{\text{Inf}}$.
  • $\theta_{\text{Pred}}$ is updated based on gradients flowing back from $L_{\text{Pred}}^{*}$ (via $\bm{h}$) and potentially $L_{\text{Inf}}$ (via $\bm{h}$, if Enc is used). The dependency of $\bm{z}^{*}$ on $\bm{h}$ and $\theta_{\text{Dec}}$ can be handled by treating $\bm{z}^{*}$ as fixed post-inference for simplicity, or through more complex methods like implicit differentiation [79] or unrolling the inference loop [39]. In this work, $\bm{z}^{*}$ is treated as fixed after its $K$-step inference for updating $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$.
The sparsity penalty $\lambda\|\bm{z}\|_{1}$ directly shapes the target $\bm{z}^{*}$ during its $K$-step inference, thereby influencing all subsequent parameter updates.
3.6 Overall Training Algorithm
The training follows a mini-batch stochastic gradient descent procedure:
  1. 1.
    Initialization: Set up data loaders, initialize network parameters $(\theta_{\text{Pred}},\theta_{\text{Dec}})$, and choose hyperparameters ($W$, $m$, $N$, $\tau_{i}$, $\lambda$, $K$, $\alpha_{\text{inf}}$, batch size $B$, optimizer learning rate $\eta$). The Encoder $\theta_{\text{Enc}}$ is optional for the core prediction task if inference is always done via K-step optimization.
  2. 2.
    Epoch Loop: For each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$:
    1. (a)
      Compute history encoding: $\bm{h}^{(j)}=\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
    2. (b)
      Infer optimal latent variable via $K$-step gradient descent:
      $\bm{z}^{*(j)}=\arg\min_{\bm{z}}E(Y^{(j)},\bm{h}^{(j)},\bm{z})$
      (using current $\theta_{\text{Dec}}$)
    3. (c)
      Calculate per-sample prediction loss: $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)};\theta_{\text{Dec}}))$
    4. (d)
      Compute average batch prediction loss: $L_{\text{Pred}}^{*}=\frac{1}{B}\sum_{j=1}^{B}L_{\text{Pred}}^{*(j)}$
    5. (e)
      Compute gradients of $L_{\text{Pred}}^{*}$ with respect to $\theta_{\text{Pred}},\theta_{\text{Dec}}$.
    6. (f)
      Update parameters $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using the chosen optimizer (e.g., Adam [54]).
  3. 3.
    Validation & Tuning: Periodically evaluate $L_{\text{Pred}}^{*}$ (using $\bm{z}^{*}$ inferred via $K$-step optimization at validation time) and sparsity metrics (e.g., $\|\bm{z}^{*}\|_{0}$) on a validation set. Use results to guide hyperparameter tuning (especially $\lambda$, $K$, $m$) and for early stopping [78].
@@ 3.6 Overall Training Algorithm
    Testing & Analysis: Evaluate the final trained model on a held-out test set using $L_{\text{Pred}}^{*}$ as the primary performance metric (with $\bm{z}^{*}$ inferred via $K$-step optimization). Analyze the inferred latent variables $\bm{z}^{*}$ from the test set—examining their sparsity, correlations with input features, and temporal evolution—to assess interpretability.
@@ 3.7 Interpretability via Sparsity
The explicit use of the L1 norm $R(\bm{z})=\|\bm{z}\|_{1}$ in the energy function plays a critical role in achieving interpretability [94]. By encouraging components of the optimal latent variable $\bm{z}^{*}$ to be exactly zero, the model is forced to represent the underlying market dynamics responsible for predicted price changes using a minimal set of active latent factors. Analyzing which factors are active can provide insights into the model’s reasoning. The hyperparameter $\lambda$ directly controls this sparsity.
@@ 4.1 Data Sources
Data was aggregated from multiple publicly accessible or commonly subscribed financial data provider sources:
  • Financial Information Providers: Services like Refinitiv Eikon or Bloomberg (via institutional access), and publicly available APIs from exchanges (e.g., CME Group, ICE) or financial data aggregators (e.g., Quandl - now part of Nasdaq Data Link, Yahoo Finance) — for daily market prices, futures data, and economic indicators.
  • Commodity Exchanges: Direct data feeds or historical data services from London Metal Exchange (LME), CME Group (NYMEX), and Intercontinental Exchange (ICE).
  • Government Agencies: U.S. Energy Information Administration (EIA), Bureau of Labor Statistics (BLS), Federal Reserve Economic Data (FRED), and National Bureau of Statistics of China.
  • International Organizations: International Monetary Fund (IMF), World Bank, Organisation for Economic Co-operation and Development (OECD).
  • Specialized Data Providers: Baker Hughes (rig counts via public releases), Commodity Futures Trading Commission (CFTC) (Commitment of Traders reports), and NOAA/NWS or equivalent national/international meteorological services (for historical weather data).
While the primary price data for model training is daily, several fundamental indicators are incorporated at lower frequencies. Table 2 illustrates sources for some representative monthly variables.
Table 2: Illustrative sources for some monthly input variables used in this research.
@@
Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. Energy Information Administration
(https://www.eia.gov/dnav/pet/hist/RBRTED.htm)
Freely available
Copper Price (benchmark) Monthly \$/ton World Bank Commodity Price Data (Pink Sheet)
(https://www.worldbank.org/en/research/commodity-markets)
Freely available
Copper Supply, Demand Indicators Monthly Various U.S. Geological Survey (USGS) Mineral Commodity Summaries
(https://www.usgs.gov/centers/national-minerals-information-center/mineral-commodity-summaries)
Freely available
Crude Oil Supply and Demand Indicators Monthly Various U.S. Energy Information Administration (EIA) Short-Term Energy Outlook (STEO) Freely available
@@ 4.2 Feature Selection
A wide array of features were selected, categorized as:
@@ 4.2.1 Target Variable (Daily Futures Prices)
  • Copper: LME Official Cash Settlement Price (or nearby futures contract).
  • Crude Oil: WTI front-month futures price (NYMEX); Brent Crude front-month futures (ICE).
@@ 4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest rates (central bank rates, sovereign yields like US 10-Year Treasury), currency rates (DXY, EUR/USD, USD/CNY).
@@ 4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories (LME, COMEX, SHFE), publicly reported global/major producer output, China’s consumption/imports, scrap indicators (where publicly accessible). Crude Oil: EIA weekly inventories, US production, publicly available OPEC+ figures, Baker Hughes rig counts, IEA/OPEC demand estimates.
@@ 4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, major equity indices (S&P 500, FTSE 100, Shanghai Composite).
@@ 4.2.6 Technical Indicators (Calculated from Daily Prices)
Moving Averages (SMA, EMA), Oscillators (RSI, MACD), ADX [70].
@@ 4.2.7 Other (Exploratory)
Inter-commodity spreads, relevant weather indices (e.g., ENSO, hurricane impact assessments if data accessible).
@@ 4.3 Data Frequency and Time Period
Primary modeling frequency is daily. Lower frequency data (weekly, monthly, quarterly) are forward-filled, strictly aligning with information release dates to prevent lookahead bias [34]. Particular care is taken with lagged official statistics (e.g., STEO) to ensure only historically available information is used at each time step $t$. The dataset spans January 1, 2005 – December 31, 2023.
@@ 4.4 Data Pre-processing
Standard steps are applied [37]:
  1. 1.
    Missing Data: Features with $\geq$40missing values removed; remaining gaps primarily forward-filled.
  2. 2.
    Feature Transformation: Log returns or percentage changes for non-stationary series. Log transformation for skewed positive features.
  3. 3.
    Normalization/Standardization: Z-score normalization, with parameters derived only from the training set.
  4. 4.
    Outlier Treatment: Applied cautiously (e.g., winsorization at 1st/99th percentiles) to preserve genuine volatility.
@@ 4.5 Final Dataset Construction
Input sequences $X_{t}\in\mathbb{R}^{W\times d}$ (look-back $W=60$ days) and targets $Y_{t}\in\mathbb{R}^{N}$ (log prices at $N=3$ horizons: $\tau_{1}=1,\tau_{2}=5,\tau_{3}=22$ days) are formed. Chronological splits: Training (2005–2018), Validation (2019–2020), Test (2021–2023) [9].
@@ 5.1 Experimental Setup
  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [89]). Dec & Enc (training only): MLPs (hidden layers 64, 32 units; ReLU). Latent dimension $m=16$.
  • Inference: $K=10$ gradient steps, $\alpha_{\text{inf}}=0.01$. At test time, inference starts from $\bm{z}^{(0)}=0$.
@@
Persistence, ARIMA (orders via ‘pmdarima‘ [2, 87]), XGBoost [20] (tuned via CV, separate model per horizon), Standard LSTM (similar to RSAE’s Pred, direct output).
@@
RMSE, MAE, Directional Accuracy (DA) on log prices [48]. Average $L_{0}$ norm for RSAE sparsity.
@@ 5.2 Predictive Accuracy Results
Observations from Accuracy Results: The RSAE framework generally achieves competitive or superior predictive performance compared to the baselines.
  • Error Metrics (RMSE, MAE): RSAE consistently yields the lowest RMSE/MAE values across horizons for both commodities, suggesting higher accuracy in predicting price change magnitudes. Improvements are more noticeable at shorter horizons.
  • Directional Accuracy (DA): RSAE also shows strong DA, often outperforming baselines. ML-based models significantly outperform Persistence and ARIMA in DA.
  • Performance Degradation: As expected, accuracy decreases for all models as the forecast horizon extends.
  • Sparsity of RSAE: The average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil) out of $m=16$, indicating effective sparsity from L1 regularization.
These results suggest RSAE’s architecture can improve forecasting accuracy while enabling interpretability.
@@ 5.3 Interpretability Analysis of Latent Factors
Analysis of $\bm{z}^{*}$ on the test set provided insights.
Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active, varying over time, suggesting adaptive representation.
Correlation with Input Features and Qualitative Economic Interpretation
Correlations between active $\bm{z}^{*}_{k}$ and key input features informed qualitative interpretations. Copper (Example Interpretations):
  • $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory): Positively correlated with Global PMI/Chinese industrial production, negatively with LME inventories. Active during economic acceleration.
  • $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD): Positively correlated with DXY/US Treasury yields. Stronger dollar/rates often showed this factor active with downward price pressure.
  • $z^{*}_{Cu,3}$ (Supply Shock Indicator): Activated during reported supply disruptions (e.g., mine strikes), often coinciding with short-term rallies.
  • $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): Correlated with VIX and COT speculative positioning.
Crude Oil (Example Interpretations):
  • $z^{*}_{Oil,1}$ (US Demand & Inventory): Correlated with EIA US inventories (negatively with builds) and mobility/gasoline demand.
  • $z^{*}_{Oil,2}$ (Geopolitical Risk & OPEC+ Policy): Spiked around OPEC+ announcements and geopolitical events.
  • $z^{*}_{Oil,3}$ (Market Risk Appetite): Correlated with credit spreads, equity shifts, and OVX.
  • $z^{*}_{Oil,4}$ (Global ex-US Demand): Correlated with global growth forecasts and non-US industrial activity.
Case Study Examples
  • Russia-Ukraine War (Early 2022): For Crude Oil, $z^{*}_{Oil,2}$ increased dramatically, aligning with the oil price rally and geopolitical risks.
  • Recession Fears/China Lockdowns (Mid-Late 2022): For Copper, $z^{*}_{Cu,1}$ turned significantly negative during high recession concerns and China’s lockdowns, coinciding with copper price weakness.
These examples illustrate how RSAE’s sparse factors can align with market narratives.
@@ 5.4 Discussion of Results
The RSAE framework achieves competitive predictive accuracy while inducing a sparse latent representation. The interpretability analysis, though qualitative, suggests learned factors capture meaningful economic influences (e.g., ”Global Demand,” ”Geopolitical Risk”), a key advantage over black-box models. RSAE navigates the accuracy-interpretability trade-off by incorporating L1 sparsity in its energy-based formulation. These findings position RSAE as a promising approach for interpretable commodity futures forecasting.
@@ 6 Discussion and Conclusion
This study introduced and evaluated a Regularized Sparse Autoencoder (RSAE) framework for multi-horizon commodity futures forecasting and interpretable factor discovery. The experimental results suggest its potential.

### v2

@@ 3.6 Overall Training Algorithm
    Initialize $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{XtoZ}}$.
@@
    For each training batch $\{(X^{(j)},Y^{(j)})\}$:
    1. (a)
      $\bm{h}^{(j)}=\text{Pred}(X^{(j)})$.
    2. (b)
      Infer $\bm{z}^{*(j)}$ via $K$-step minimization of $E(Y^{(j)},\bm{h}^{(j)},\bm{z})$.
    3. (c)
      $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)}))$.
    4. (d)
      Update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ via gradients of $\text{mean}(L_{\text{Pred}}^{*})$ (e.g., Adam [56]).
    5. (e)
      Update $\theta_{\text{XtoZ}}$ to predict $\bm{z}^{*(j)}$ from $X^{(j)}$.
@@
    Validate using $L_{\text{Pred}}^{*}$ (with $\bm{z}_{\text{val}}=\text{Enc}_{\text{XtoZ}}(X_{\text{val}})$) for tuning and early stopping [82].
@@
    Test similarly. Analyze inferred $\bm{z}_{\text{test}}$ for interpretability.
@@ 3.7 Interpretability via Sparsity
The L1 norm encourages sparse $\bm{z}^{*}$, aiding interpretability [97].
@@ 4.1 Data Sources
Aggregated from public/common sources: Refinitiv Eikon, Bloomberg, exchange APIs (CME, ICE, LME), government agencies (EIA, BLS, FRED, NBS China), international organizations (IMF, World Bank, OECD), and specialized providers (Baker Hughes, CFTC, NOAA/NWS). Table 2 gives examples.
Table 2: Illustrative sources for some monthly input variables.
@@
Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. EIA (Pet. Hist.) Freely available
Copper Price (benchmark) Monthly \$/ton World Bank (Comm. Data) Freely available
Copper Supply, Demand Ind. Monthly Various USGS (Min. Sum.) Freely available
Crude Oil Supply, Demand Ind. Monthly Various U.S. EIA Short-Term Energy Outlook (STEO) Freely available
@@ 4.2.1 Target Variable (Daily Futures Prices)
Copper (LME nearby futures), Crude Oil (WTI, Brent front-month).
@@ 4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest/currency rates.
@@ 4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories, producer output, China consumption/imports, scrap indicators. Crude Oil: EIA weekly inventories, US/OPEC+ production, rig counts, demand estimates.
@@ 4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Comp.
@@ 4.2.6 Technical Indicators (Calculated from Daily Prices)
SMAs, EMAs, RSI, MACD, ADX [72].
@@ 4.2.7 Other (Exploratory)
Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
@@ 4.3 Data Frequency and Time Period
Primary modeling is daily. Lower frequency data forward-filled, with release lags respected for official stats (e.g., STEO) to prevent lookahead bias [35]. Dataset: Jan 1, 2005 – Dec 31, 2023.
@@ 4.4 Data Pre-processing
Standard steps [38]: Missing data handling (remove if $\geq$40%, else ffill), transformations for stationarity/skewness (log returns, log transform), Z-score normalization (params from training set only), cautious outlier treatment.
@@ 4.5 Final Dataset Construction
Inputs $X_{t}\in\mathbb{R}^{W\times d}$ ($W=60$), targets $Y_{t}\in\mathbb{R}^{N}$ ($N=3$ horizons: 1, 5, 22 days). Chronological splits: Train (2005–2018), Valid (2019–2020), Test (2021–2023) [9]. Robustness could be improved with rolling-origin CV [10].
@@ 5 Experimental Setup and Results
The empirical results presented herein are based on the models trained and evaluated according to the setup detailed below.
@@ 5.1 Experimental Setup
  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [92]). Dec: MLP (hidden layers 64, 32 units; ReLU). A separate encoder $\text{Enc}_{\text{XtoZ}}(X)$ maps $X$ to $\bm{z}^{*}$ at test time. Latent dimension $m=16$.
  • Inference (Training): $K=10$ gradient steps for $\bm{z}^{*}$, $\alpha_{\text{inf}}=0.01$.
@@
Persistence, ARIMA [14] (orders via ‘pmdarima‘ [2, 90]), XGBoost [21] (tuned via CV), Standard LSTM. Baseline hyperparameter tuning was standard; a fully balanced budget vs. RSAE is a future work consideration.
@@
RMSE, MAE, Directional Accuracy (DA) on log prices [50]. Average $L_{0}$ norm for RSAE. Days with zero price change are included; if prediction is also no change, it’s a correct DA. Advanced metrics like Pesaran-Timmermann test [79] are for future work.
$DA_{\tau_{i}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
@@ 5.2 Predictive Accuracy Results
Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance. RMSE/MAE values are consistently lowest. DA is strong. Performance degrades with horizon. Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
@@ 5.3 Interpretability Analysis of Latent Factors
5.3.1 Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active.
5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
The interpretations are qualitative, based on correlations and event alignments. See Appendix A for supporting correlation details. Copper (Examples): $z^{*}_{Cu,1}$ (Global Demand/Inventory), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows). Crude Oil (Examples): $z^{*}_{Oil,1}$ (US Demand/Inventory), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
5.3.3 Case Study Examples
Alignments with Russia-Ukraine War and China Lockdowns support factor interpretations.
@@ 5.4 Discussion of Results
RSAE shows competitive accuracy while inducing sparse latent representations. Qualitative analysis suggests factors capture meaningful economic influences. The L1 sparsity aids the accuracy-interpretability balance. Results are from a specific run; further validation (seeds, data perturbations, statistical significance tests like Diebold-Mariano) is needed.
@@ 6 Discussion and Conclusion
This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery.